# Efficient Learned Wavelet Image and Video Coding

### v1

@@ 3.1 Image Compression Model: pWave++
(b) Remaining subband types $\tilde{\bm{s}}_{i},i=2,\ldots,12$
@@
Our pWave++ model follows the structure of the end-to-end wavelet image coding framework iWave++. Fig. 1 provides an overview of the approach. The input $\bm{x}$ is one color component of an image in the YCbCr color space. The analysis transform corresponds to a learned discrete wavelet transform (DWT) and the synthesis transform to the inverse discrete wavelet transform (IDWT). The latent space $\bm{y}$ hence has a known structure and represents a wavelet decomposition. The subband layout is shown in the middle of Fig. 1 with two decomposition levels for visualization. Our model has four decomposition levels in total. The subbands $\bm{y}$ are quantized by scalar quantization according to $\tilde{\bm{y}}=\lfloor\bm{y}\cdot\Delta\rceil$, where straight-through gradients are used for the rounding operation during training. At the decoder side, the reconstructed subbands are obtained as $\hat{\bm{y}}=\tilde{\bm{y}}/\Delta$. We use a dedicated trainable parameter $\Delta_{\mathrm{L}}$ for the lowpass subband and another parameter $\Delta_{\mathrm{H}}$ for the remaining subband types, in contrast to the baseline model with a single parameter for all subbands. As indicated by the red arrows in Fig. 1, the quantized subbands $\tilde{\bm{y}}$ are coded sequentially. The model uses adaptive arithmetic coding to compress the quantized wavelet representation. The context model for entropy parameter estimation will be explained in the following Section 3.2. We employ a Laplacian distribution as probability model. After the inverse wavelet transform, the final reconstruction $\hat{\bm{x}}$ is obtained from a post-processing module. Because the encoding and decoding transform share their parameters, this module is required to compensate quantization artifacts.
@@ 4.2 Image Coding: pWave++
Our pWave++ model achieves similar rate-distortion performance compared to the baseline iWave++ model, as shown in Fig. 7. We measure a decreased coding efficiency with an increased Bjøntegaard delta (BD) bitrate of +1.5 % in comparison to the baseline model. Note that this deterioration is caused by the lowest rate point of pWave++ alone, as its rate-distortion curve otherwise matches the curve of iWave++ (cf. Fig. 7). Table 1 provides runtime measurements comparing the baseline model iWave++ to our proposed pWave++. The runtimes are averaged over all images from the Kodak data set and performed on a NVIDIA RTX A6000 GPU. Both the encoding and decoding time significantly decrease with our pWave++ model. We measure around 6 seconds decoding time instead of 1995 seconds, which corresponds to a speedup factor of 353.
@@
Using the autoregressive context fusion model for the first subband to be coded, i.e., the LL subband, is important for the rate-distortion performance of pWave++. We demonstrate this by training a pWave++ model with a four-step context fusion model for the LL subband. As can be seen in Table 1, the full four-step model achieves an even faster decoding time. Omitting the autoregressive model leads to faster decoding by a factor of 16. However, the rate-distortion performance of the full four-step model is significantly worse, as can be seen from the rate-distortion curves in Fig. 7. This behavior is caused by a different subband layout of the full four-step pWave++ model: Coding the LL subband without spatial context lets the model store a lowpass representation in the subbands LH1, HL1, and HH1, and creates a highpass representation towards the lower decomposition levels, with the sparsest subbands being LL4, LH4, HL4, and HH4. This finding highlights that the subband layout known from the traditional wavelet transform - with a lowpass representation in the LL4 subband and with sparse directional details in the remaining subbands - is crucial for the performance of iWave++ as well as pWave++.
@@ 5 Conclusion
In this paper, we introduced a novel parallelized context model for the image coder iWave++ that achieves a decoding speedup by a factor of over 350 for image coding and of over 350 for video coding. Our experimental results demonstrated that the efficient model can be employed for wavelet image and video coding with small BD rate losses of up to 1.5%. We showed that the interpretable subband layout known from traditional wavelet transforms performs best and provided a visual analysis in comparison to other transforms in learned compression. In future work, we will focus on improving the prediction capabilities of our pMCTF model by investigating advanced methods for motion estimation and compensation or integrating multiple reference frames in MCTF.

### v2

@@ 3.1 Image Compression Model: pWave++
(b) Remaining subband types $\tilde{\bm{s}}_{i},i=2,\ldots,13$
@@
Our pWave++ model follows the structure of the end-to-end wavelet image coding framework iWave++. Fig. 1 provides an overview of the approach. The input $\bm{x}$ is one color component of an image in the YCbCr color space. The analysis transform corresponds to a learned discrete wavelet transform (DWT) and the synthesis transform to the inverse discrete wavelet transform (IDWT). The latent space $\bm{y}$ hence has a known structure and represents a wavelet decomposition. The subband layout is shown in the middle of Fig. 1 with two decomposition levels for visualization. Our model has four decomposition levels in total. The subbands $\bm{y}$ are quantized by scalar quantization according to $\tilde{\bm{y}}=\lfloor\bm{y}\cdot\Delta\rceil$, where straight-through gradients are used for the rounding operation during training. At the decoder side, the reconstructed subbands are obtained as $\hat{\bm{y}}=\tilde{\bm{y}}/\Delta$. We use a dedicated trainable parameter $\Delta_{\mathrm{L}}$ for the lowpass subband and another parameter $\Delta_{\mathrm{H}}$ for the remaining subband types, in contrast to the baseline model with a single parameter for all subbands. As indicated by the red arrows in Fig. 1, the quantized subbands $\tilde{\bm{y}}$ are coded sequentially. The model uses adaptive arithmetic coding to compress the quantized wavelet representation. The context model for entropy parameter estimation will be explained in the following Section 3.2. We employ a Laplacian distribution as probability model following [16]. After the inverse wavelet transform, the final reconstruction $\hat{\bm{x}}$ is obtained from a post-processing module. Because the encoding and decoding transform share their parameters, this module is required to compensate quantization artifacts.
@@ 4.2 Image Coding: pWave++
Our pWave++ model achieves similar rate-distortion performance compared to the baseline iWave++ model, as shown in Fig. 7. We measure a decreased coding efficiency with an increased Bjøntegaard delta (BD) bitrate of +1.5 % in comparison to the baseline model. Note that this deterioration is caused by the lowest rate point of pWave++ alone, as its rate-distortion curve otherwise matches the curve of iWave++ (cf. Fig. 7). Table 1 provides runtime measurements comparing the baseline model iWave++ to our proposed pWave++. The runtimes are averaged over all images from the Kodak data set and performed on a NVIDIA RTX A6000 GPU. Both the encoding and decoding time significantly decrease with our pWave++ model. We measure around 6 seconds decoding time instead of 1995 seconds, which corresponds to a speedup factor of 354.
@@
Using the autoregressive context fusion model for the first subband to be coded, i.e., the subband LL4, is important for the rate-distortion performance of pWave++. We demonstrate this by training a pWave++ model with a four-step context fusion model for the LL4 subband. As can be seen in Table 1, the full four-step model achieves an even faster decoding time. Omitting the autoregressive model leads to faster decoding by a factor of 16. However, the rate-distortion performance of the full four-step model is significantly worse, as can be seen from the rate-distortion curves in Fig. 7. This behavior is caused by a different subband layout of the full four-step pWave++ model: Coding the LL4 subband without autoregressive spatial context lets the model store a lowpass representation in the subbands LH1, HL1, and HH1, and creates a highpass representation towards the lower decomposition levels, with the sparsest subbands being LL4, LH4, HL4, and HH4. This finding highlights that the subband layout known from the traditional wavelet transform - with a lowpass representation in the LL4 subband and with sparse directional details in the remaining subbands - is crucial for the performance of iWave++ as well as pWave++.
@@ 5 Conclusion
In this paper, we introduced a novel parallelized context model for the image coder iWave++ that achieves a decoding speedup by a factor of over 350 for image coding and of over 240 for video coding. Our experimental results demonstrated that the efficient model can be employed for wavelet image and video coding with small BD rate losses of up to 1.5%. We showed that the interpretable subband layout known from traditional wavelet transforms performs best and provided a visual analysis in comparison to other transforms in learned compression. In future work, we will focus on improving the prediction capabilities of our pMCTF model by investigating advanced methods for motion estimation and compensation or integrating multiple reference frames in MCTF.